OpenStudy (anonymous):

Solve: y^2 + 8y + 6 = 0

OpenStudy (anonymous):

hi use the quadratic formula on y^2 + 8y + 6 = 0 y=[-8-+sqrt(8)^2 -4(1)(6)]/2(1) y=[-8-+sqrt(64) -24]/2(1) y=[-8-+sqrt40]/2(1) y=[-8-+2sqrt10]/2(1) y=[-4-+sqrt10] y=-4-sqrt10 ans y=-4+sqrt10 ans